OpenStudy (anonymous):

which of the point slope equations below is correct for the line that passes through points (6,5) and (3,3)? check all that apply

OpenStudy (anonymous):

A.y-3=3/2(x-3) B. y+5=2/3(x+6) c. y-5=2/3(x+6) D. y-3=2/3(x+3) E. y+6=3/2(x-5) F. y-3=2/3(x-3)

OpenStudy (anonymous):

@kirbykirby

OpenStudy (kirbykirby):

WHat you should do is plug in in coordinate values into the equation (like for (6,5) , plug in y = 6 and x = 6... ) The ones that work are the ones in which both sides of the equality will be equal to each other

OpenStudy (anonymous):

so the are the answers

OpenStudy (kirbykirby):

y = 6 x = 5* minor typo -_-

OpenStudy (anonymous):

I don't kno the answers

OpenStudy (anonymous):

flutter youuuu dude :)

OpenStudy (anonymous):

f u c k you dude

OpenStudy (kirbykirby):

o_o